# A multi-band study of pulsar glitches with Fermi-LAT and Parkes

### v1

@@ 6 CONCLUSIONS
We connected the observed glitches with the dynamics of neutron vortex arrays inside the pulsars and discussed the theoretical interpretation of these events in the framework of vortex creep model. For our sample of glitches, application of the vortex creep model revealed that the total fractional moment of inertia of the crustal superfluid required to tapped at glitches is $\lesssim 3.5\%$, indicating that only crustal superfluid and probably some part of the outer core region of a neutron star are involved in glitches. Yet another implication of the application of the vortex creep model is that the model estimate for the time to the next glitch in general agrees well with the observed inter-glitch time for a given pulsar’s particular glitch. Thus, by doing timing fit to a particular glitch future observations can be planned for corresponding pulsar to establish any variations in the pulse profile concurrent with the glitch occurrence. Such a detection will enable us to connect the magnetospheric physics to the interior dynamics and disclose the interplay between them. Different parts of the post-glitch recovery, in particular the exponential relaxation can be studied to infer microphysics related to the neutron star crustal and core matter. We plan to compute the relaxation time derived from microscopic two- and three-body nuclear forces, incorporating pasta phases and proton effective masses self-consistently, as well as application to the observed cases in a separate paper. We’d also like to address the effects of more updated superfluidity/superconductivity-related interaction parameters by taking proximity etc. into account on the exponential recovery time estimate. Simulating pulsar glitches through experimental observations in rotating ultracold atoms offers exciting potential to constrain the complex physics associated with superfluid vortices (Poli et al., 2023). Future astrophysical observations will provide a larger sample, particularly with variations in post-glitch behaviour, promising greater precision in glitch constraints and offering insights into the dynamics of superfluid vortices beyond the coarse-grained two-component scenario. Additionally, pulsar glitches may imprint themselves on gravitational wave radiation, establishing a promising link between multi-messenger studies of these phenomena and insights into the dense matter EOS (Cheunchitra et al., 2024; Wilson & Ho, 2024; Yim et al., 2024).
@@ Acknowledgements
The work is supported by National SKA Program of China (no. 2020SKA0120300), the Strategic Priority Research Program of the Chinese Academy of Sciences (grant no. XDB0550300), the National Natural Science Foundation of China (grant nos 12273028, 12041304 and 12288102), and the Major Science and Technology Program of Xinjiang Uygur Autonomous Region (grant no. 2022A03013, 2022A03013-4). EG acknowledges support from the National Natural Science Foundation of China (NSFC) programme 11988101 under the foreign talents grant QN2023061004L. ZZ is supported by the Natural Science Basic Research Program of Shaanxi (program no. 2024JC-YBQN-0036). SD is supported by the Guizhou Provincial Science and Technology Foundation (no. ZK[2022]304) and the Scientific Research Project of the Guizhou Provincial Education (no. KY[2022]132). PW acknowledges support from the National Natural Science Foundation of China under grant U2031117 and the Youth Innovation Promotion Association CAS (id. 2021055). DL is supported by the 2020 project of Xinjiang Uygur autonomous region of China for flexibly fetching in upscale talents. We acknowledge the use of the public data from the Fermi-LAT data archive. The Parkes radio telescope is part of the Australia Telescope National Facility which is funded by the Commonwealth of Australia for operation as a National Facility managed by CSIRO. This paper includes archived data obtained through the CSIRO Data Access Portal.

### v2

@@ 6 CONCLUSIONS
We connected the observed glitches with the dynamics of neutron vortex arrays inside the pulsars and discussed the theoretical interpretation of these events in the framework of vortex creep model. For our sample of glitches, application of the vortex creep model revealed that the total fractional moment of inertia of the crustal superfluid required to tapped at glitches is $\lesssim 3.5\%$, indicating that only crustal superfluid and probably some part of the outer core region of a neutron star are involved in glitches. Yet another implication of the application of the vortex creep model is that the model estimate for the time to the next glitch in general agrees well with the observed inter-glitch time for a given pulsar’s particular glitch. Thus, by doing timing fit to a particular glitch future observations can be planned for corresponding pulsar to establish any variations in the pulse profile concurrent with the glitch occurrence. Such a detection will enable us to connect the magnetospheric physics to the interior dynamics and disclose the interplay between them. Different parts of the post-glitch recovery, in particular the exponential relaxation can be studied to infer microphysics related to the neutron star crustal and core matter. We plan to compute the relaxation time derived from microscopic two- and three-body nuclear forces, incorporating pasta phases and proton effective masses self-consistently, as well as application to the observed cases in a separate paper (Tu & Li, 2024). We’d also like to address the effects of more updated superfluidity/superconductivity-related interaction parameters by taking proximity etc. into account on the exponential recovery time estimate. Simulating pulsar glitches through experimental observations in rotating ultracold atoms offers exciting potential to constrain the complex physics associated with superfluid vortices (Poli et al., 2023). Future astrophysical observations will provide a larger sample, particularly with variations in post-glitch behaviour, promising greater precision in glitch constraints and offering insights into the dynamics of superfluid vortices beyond the coarse-grained two-component scenario. Additionally, pulsar glitches may imprint themselves on gravitational wave radiation, establishing a promising link between multi-messenger studies of these phenomena and insights into the dense matter EOS (Cheunchitra et al., 2024; Wilson & Ho, 2024; Yim et al., 2024).
@@ Acknowledgements
This work is supported by National SKA Program of China (no. 2020SKA0120300), the Strategic Priority Research Program of the Chinese Academy of Sciences (grant no. XDB0550300), the National Natural Science Foundation of China (grant nos 12494572, 12273028, 12041304 and 12288102), and the Major Science and Technology Program of Xinjiang Uygur Autonomous Region (grant no. 2022A03013, 2022A03013-4). EG acknowledges support from the National Natural Science Foundation of China (NSFC) programme 11988101 under the foreign talents grant QN2023061004L. ZZ is supported by the Natural Science Basic Research Program of Shaanxi (program no. 2024JC-YBQN-0036). SD is supported by the Guizhou Provincial Science and Technology Foundation (no. ZK[2022]304) and the Scientific Research Project of the Guizhou Provincial Education (no. KY[2022]132). PW acknowledges support from the National Natural Science Foundation of China under grant U2031117 and the Youth Innovation Promotion Association CAS (id. 2021055). DL is supported by the 2020 project of Xinjiang Uygur autonomous region of China for flexibly fetching in upscale talents. We acknowledge the use of the public data from the Fermi-LAT data archive. The Parkes radio telescope is part of the Australia Telescope National Facility which is funded by the Commonwealth of Australia for operation as a National Facility managed by CSIRO. This paper includes archived data obtained through the CSIRO Data Access Portal.